# The Early Career Workshop of GR–Amaldi 2025

### v1

@@ 1 Youthful science: The need for an Early Career Workshop
Although the modern gravitational science, incorporating general relativity and gravitational waves, has been established for over 100 years, recent observations sparked the rapid growth in its community. The first direct observation of GWs [1] not only verified directly Einstein’s predictions [2, 3], but also highlighted the promise of many exciting discoveries in the years to come [4, 5, 6]. At the same time, advances in computing and the rise of big data allowed for conceptually new approaches in analysing the data and modelling complicated cosmic systems [7, 8, 9, 10, 11, 12]. This has brought many young scientists to the field of gravity, who shared not only the passion for science, but also the desire to develop their skills for their future careers. This field has become a unique ground for uniting scientists across a broad range of topics and building a strong community for the upcoming decades of scientific breakthroughs.
@@
[…]
@@ 1.1 Motivating opportunity
The International Conference on General Relativity and Gravitation (GR) is held every three years and is the principal international meeting for scientists working in gravity and relativity. The Edoardo Amaldi Conference on Gravitational Waves (Amaldi) is held every two years and is the principal international meeting for scientists working in GW science. Every six years, the GR and Amaldi meetings are held together; this joint meeting brings together experts from across classical and quantum gravity, mathematical and applied relativity, GW instrumentation and data analysis, multimessenger astronomy, relativistic astrophysics and cosmology, as well as education, outreach and community building in support of these areas. Conference attendance is an important part of academic career development [36, 37]. The gap between joint GRAmaldi meetings makes them a rare opportunity for ECRs: most PhD students may only attend one during the course of their doctoral studies. Given the significance of the joint meeting, and the lack of in-person meetings over the last few years, the GRAmaldi meeting in Glasgow was identified as an important opportunity to bring together ECRs in the field.
The size of the GRAmaldi meeting means that a large number of ECRs are expected to attend. At the meeting, many ECRs would present their work, attend presentations on topics outside of their main area, and interact with senior figures in the field. This makes the meeting an ideal setting for an Early Career Workshop while highlighting its benefits: Technical sessions would help orient attendees ahead of the diverse program of the main meeting, transferable-skills sessions would prepare attendees to make the most out of the main meeting, and networking opportunities would help them build connections to avoid feeling lost in the crowd of hundreds of attendees of the main meeting.
@@ 2 Workshop programme
The Workshop spanned three days. The first two days featured the core programme: a mix of talks and panel discussions. As with the previous GWECS workshop [39], the schedule was split roughly evenly between technical topics and transferable skills. These sessions were deliberately interleaved, so participants could still engage with a balanced programme even if they only attended part of it. The third day was dedicated to networking and social activities, aimed at consolidating the connections built during the earlier sessions. The complete timetable is given in Table 1.222On Sunday 13 July (the day between the Workshop and the main GRAmaldi meeting) several satellite events were held. ECRs were invited to a Wikipedia Edit-a-thon, which provided training on editing Wikipedia and included collaborative work on updating pages relevant to the meeting. In the evening, a public lecture was delivered by Dr. Erin Macdonald, one of the Workshop panellists. Her talk, Spacetime in Star Trek, combined reflections on her career with a discussion of how scientific concepts are represented in popular media.
@@ 2.2 Lectures
The opening lecture, by Scott Hughes, introduced black hole spacetimes, covering their unique theoretical properties [44, 45, 46], observables [47, 48, 25], modelling techniques [49, 50], and recent GW measurements [1, 3]. Katy Clough followed with a worked-example approach to numerically solving Einstein’s equations, highlighting pitfalls from ignored biases or assumptions [51, 52]. Both lectures prompted wide-ranging questions spanning theory and practical methods. Jen Gupta discussed her path into science communication, showing how outreach impacts public perception and how innovative tools like 3D printing can expand accessibility [53, 54]. Mikhail Korobko then offered strategies for navigating large conferences, from goal-setting to mental health breaks and effective follow-ups. Christopher Berry’s session on academic writing outlined techniques for structuring papers, keeping text concise and precise [55], and streamlining with macros. The day closed with Bianca Dittrich on quantum gravity approaches at the Planck scale [56], and John Veitch on GW data analysis [57, 58], demonstrating Bayesian methods, population studies, and open challenges for next-generation detectors.
Day 2 maintained a similar balance between different topics. Antonio Carillo reviewed electromagnetic signatures of relativistic systems [59, 60, 61], while Andrew Taylor connected cosmological surveys to gravitational science [62, 63, 64, 65], highlighting how future observatories could challenge the standard model [66]. Daniela Doneva argued for modified gravity in the strong-field regime [67, 68, 69], followed by Elisa Maggio on observational tests of general relativity [2, 3] and pitfalls of false deviations from it [70, 71]. Jonathan Gair concluded with an overview of GW experiments across the spectrum [4, 72, 73, 74, 16, 75, 76, 77, 78, 79], stressing the opportunities (and challenges) posed by expanding data and new detector regimes.
2.3 Panels
@@
The Careers in Academia panel combined early-career and senior researchers. Anna Green (Maastricht University) and Simone Mastrogiovanni (Istituto Nazionale di Fisica Nucleare, Rome) had both recently taken up faculty roles, while Christiana Pantelidou (University College Dublin) and Adam Pound (University of Southampton) contributed experience from research fellowships and long-term posts. Discussion ranged from career trajectories and work–life balance to the legal and personal challenges of short-term contracts.
The Careers Outside Academia panel was shaped by practical constraints: in the absence of travel funding, the panellists were primarily invited from University of Glasgow graduates. Attendance required either a supportive employer or taking time off, and in either case, this was easier to justify when the time away from work was minimized, and so was easier for those working near Glasgow. Scheduling the panel at the end of the day on Friday minimised the time away from work needed to attend. Despite the constraints, the line-up was diverse, spanning law, finance, industry, and media. Speakers included Rebecca Douglas (Patent Attorney, Hindles), Peter Wakeford (R&D Engineer, Optos), Jennifer Toher (Senior Architect, JPMorgan Chase), and Erin Macdonald (science advisor to the Star Trek franchise). Sebastian Khan (Data Scientist, Starling Bank), a Cardiff University alumnus, spoke on continuing GW-related research at their current role. Discussion focused on career paths, milestones, and timing of transitions. Panellists also addressed how far doctoral research influenced their current work and what additional training proved essential as well as work–life balance with their roles. A recurring theme was the value of the skills provided by a PhD, and the value of networking.
@@ 3 Reflections
Overall, the main theme of the responses was the good balance between the transferable-skill sessions and science introductions, as well as the chance to meet the peers. This matched well with our goals of the conference, and highlighted the need of such events within the communities.
@@ 4 Conclusions
The Workshop demonstrated the importance of investing in structured opportunities for ECRs to build community, develop cross-disciplinary understanding, and acquire skills that extend beyond research. In a field that is rapidly expanding and increasingly collaborative, such initiatives strengthen both the individuals involved and the community as a whole. Future GRAmaldi joint meetings should be encouraged to continue to support and expand similar events, ensuring that the next generation of researchers is well prepared to meet the scientific and professional challenges of gravitational science.

### v2

@@ 1 Youthful science: The need for an Early Career Workshop
Although modern gravitational science, incorporating general relativity and gravitational waves, has been established for over 100 years, recent observations have sparked a rapid growth in its community. The first direct observation of GWs [6] not only verified directly Einstein’s predictions [7, 10], but also highlighted the promise of many exciting discoveries in the years to come [2, 11, 4]. At the same time, advances in computing and the rise of big data allowed for conceptually new approaches in analysing the data and modelling complicated cosmic systems [56, 79, 23, 20, 47, 46]. This has brought many young scientists to the field of gravity, who share not only the passion for science, but also the desire to develop skills for their future careers. This field has become a unique ground for uniting scientists across a broad range of topics and building a strong community for the upcoming decades of scientific breakthroughs.
@@
The Gravitational-Wave Early Career Scientists (GWECS) initiative was established to address exactly these needs: to provide ECRs with training, community, and visibility within large international collaborations. [^1] GWECS webpage gwecs.org. Building a career in gravitational physics today almost inevitably involves working in large, distributed teams, navigating authorship and credit in publications with thousands of co-authors, and developing transferable skills in project management, communication and collaboration. These are challenges not unique to gravitational physics [73, 32], but particularly acute in this field, given the scale of current and future projects. In the context of the increasingly competitive academic environment, with limited permanent positions and growing reliance on international mobility, the demands placed upon ECRs [57, 24] makes it critical to support them through structured opportunities such as workshops. Such initiatives not only strengthen the individual careers of young scientists but also reinforce the long-term health and inclusivity of the research community as a whole.
@@ 1.1 Motivating opportunity
The International Conference on General Relativity and Gravitation (GR) is held every three years and is the principal international meeting for scientists working in gravity and relativity. The Edoardo Amaldi Conference on Gravitational Waves (Amaldi) is held every two years and is the principal international meeting for scientists working in GW science. Every six years, the GR and Amaldi meetings are held together; this joint meeting brings together experts from across classical and quantum gravity, mathematical and applied relativity, GW instrumentation and data analysis, multi-messenger astronomy, relativistic astrophysics and cosmology, as well as education, outreach and community building in support of these areas. Conference attendance is an important part of academic career development [71, 16]. The gap between joint GRAmaldi meetings makes them a rare opportunity for ECRs: most PhD students may only attend one during the course of their doctoral studies. Given the significance of the joint meeting, and the lack of in-person meetings over the last few years, the GRAmaldi meeting in Glasgow was identified as an important opportunity to bring together ECRs in the field.
The size of the GRAmaldi meeting means that a large number of ECRs were expected to attend. At the meeting, many ECRs would present their work, attend presentations on topics outside of their main area, and interact with senior figures in the field. This makes the meeting an ideal setting for an Early Career Workshop as well as highlighting the benefits of a Workshop: Technical sessions would help orient attendees ahead of the diverse program of the main meeting, transferable-skills sessions would prepare attendees to make the most out of the main meeting, and networking opportunities would help them build connections to avoid feeling lost in the crowd of hundreds of attendees of the main meeting.
@@ 2 Workshop programme
The Workshop spanned three days. The first two days featured the core programme: a mix of talks and panel discussions. As with the previous GWECS workshop [34], the schedule was split roughly evenly between technical topics and transferable skills. These sessions were deliberately interleaved, so participants could still engage with a balanced programme even if they only attended part of it. The third day was dedicated to networking and social activities, aimed at consolidating the connections built during the earlier sessions. The complete timetable is given in Table 1.222On Sunday 13 July (between the Workshop and the main meeting) there were several satellite events. Attendees were invited to a Wikipedia Edit-a-thon, which provided training on editing Wikipedia and included collaborative work updating pages relevant to the meeting. In the evening, Erin Macdonald (a Workshop panellist) delivered a public lecture, Spacetime in Star Trek, combining reflections on her career with discussion of how scientific concepts are represented in popular media.
@@ 2.2 Lectures
The opening lecture, by Scott Hughes (MIT), introduced black hole spacetimes, covering their unique theoretical properties [55, 40, 70], observables [52, 13, 19], modelling techniques [35, 41], and recent GW measurements [6, 10]. Katy Clough (Queen Mary University of London) followed with a worked-example approach to numerically solving Einstein’s equations, highlighting pitfalls from ignored biases or assumptions [33, 29]. Both lectures prompted wide-ranging questions spanning theory and practical methods. Jen Gupta (University of Portsmouth) discussed her path into science communication, showing how outreach impacts public perception and how innovative tools like 3D printing can expand accessibility [38, 65]. Mikhail Korobko (University of Hamburg) then offered strategies for navigating large conferences, from goal-setting to mental health breaks and effective follow-ups. Christopher Berry’s (University of Glasgow) session on academic writing outlined techniques for structuring papers, keeping text concise and precise [74], and streamlining with macros. The day closed with Bianca Dittrich (Perimeter Institute and University of Waterloo) on quantum gravity approaches at the Planck scale [39], and John Veitch (University of Glasgow) on GW data analysis [42, 3], demonstrating Bayesian methods, population studies, and open challenges for next-generation detectors.
Day 2 maintained a similar balance between different topics. Antonio Carillo (University College Dublin) reviewed electromagnetic signatures of relativistic systems [64, 76, 51], while Andrew Taylor (University of Edinburgh) connected cosmological surveys to gravitational science [49, 12, 43, 15], highlighting how future observatories could challenge the standard model [37]. Daniela Doneva (Universität Tübingen) argued for modified gravity in the strong-field regime [36, 50, 78], followed by Elisa Maggio (Max Planck Institute for Gravitational Physics) on observational tests of general relativity [7, 10] and pitfalls of false deviations from it [54, 60]. Jonathan Gair (Max Planck Institute for Gravitational Physics) concluded with an overview of GW experiments across the spectrum [2, 9, 58, 5, 45, 72, 31, 62, 61, 75], stressing the opportunities (and challenges) posed by expanding data and new detector regimes.
2.3 Career panels
@@
The Academia panel combined early-career and senior researchers. Anna Green (Maastricht University) and Simone Mastrogiovanni (Istituto Nazionale di Fisica Nucleare, Rome) had both recently taken up faculty roles, while Christiana Pantelidou (University College Dublin) and Adam Pound (University of Southampton) contributed experience from research fellowships and long-term posts. Discussion ranged from career trajectories and work–life balance to the legal and personal challenges of short-term contracts.
The Outside Academia panel was shaped by practical constraints: in the absence of travel funding, the panellists were primarily invited from University of Glasgow graduates. Attendance required either a supportive employer or taking time off, and in either case, this was easier to justify when the time away from work was minimized, and so was easier for those working near Glasgow. Scheduling the panel at the end of the day on Friday minimised the time away from work needed to attend. Despite the constraints, the line-up was diverse, spanning law, finance, industry, and media. Speakers included Rebecca Douglas (Patent Attorney, Hindles), Peter Wakeford (R&D Engineer, Optos), Jennifer Toher (Senior Architect, JPMorgan Chase), and Erin Macdonald (science advisor to the Star Trek franchise). Sebastian Khan (Data Scientist, Starling Bank), a Cardiff University alumnus, spoke on continuing GW-related research at their current role. Discussion focused on career paths, milestones, and timing of transitions. Panellists also addressed how far doctoral research influenced their current work and what additional training proved essential as well as work–life balance with their roles. A recurring theme was the value of the skills provided by a PhD, and the value of networking.
@@ 3 Reflections
Overall, the main theme of the responses was the good balance between the transferable-skill sessions and science introductions, as well as the chance to meet peers. This matched well with our goals for the Workshop, and highlighted the need for such events within the communities.
@@ 4 Conclusions
The Workshop demonstrated the importance of investing in structured opportunities for ECRs to build community, develop cross-disciplinary understanding, and acquire skills that extend beyond research. In a field that is rapidly expanding and increasingly collaborative, such initiatives strengthen both the individuals involved and the community as a whole. Thanks to early engagement with potential sponsors, the Workshop was organised without placing additional financial burden on attendees. Future GRAmaldi joint meetings should be encouraged to continue to support and expand similar events, ensuring that the next generation of researchers is well prepared to meet the scientific and professional challenges of gravitational science.